I need to find the interest rate of a total loan of $1600.00 to be repaid over 3 years. if the original loan amount was $1000.00. Every time I attempt to find the rate, I get it messed up.
Answer by jim_thompson5910(35256)

I'm going to assume this is a simple interest loan (and not a compound interest loan)
The original amount is $1,000
The amount paid back is $1,600
The difference is 600 dollars (1600-1000 = 600)
This difference of $600 is the total amount of interest paid.
So i = 600 where i = interest

We'll use the formula
i = P*r*t
where,
i = 600, the interest found above
P = 1000, amount loaned out (Principal)
r = unknown
t = 3 years

So we get...
i = P*r*t
600 = 1000*r*3
600 = 1000*3*r
600 = 3000*r
600/3000 = 3000*r/3000
0.2 = r
r = 0.2

Since r = 0.2, this means that the interest rate is 20% because 0.2 = 20/100 = 20%